OpenStudy (anonymous):

h(x) = 2ln((x-1)/2)+3 Range and domain this function What is the inverse of the function :)

OpenStudy (anonymous):

any takers please

OpenStudy (anonymous):

ln((x-1)/2)

OpenStudy (anonymous):

if the input is \(\frac{x-1}{2}\) then solve \[\frac{x-1}{2}>0\] for \(x\)

OpenStudy (anonymous):

the range of the log is all real numbers

OpenStudy (anonymous):

so x must be greater than 1 and y is an element of all the reals?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you want the inverse solve \[x=2\ln(\frac{y-1}{2})+3\] for \(y\) in a few steps

OpenStudy (anonymous):

only one step is not algebra subtract 3 divide by 2 not algebra step, write in exponential form etc

OpenStudy (anonymous):

I have got to \(-3/2 = ln(x-1/2)\)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i just get so confused with exponentials etc :(

OpenStudy (anonymous):

k lets go slow

OpenStudy (anonymous):

\[x=2\ln(\frac{y-1}{2})+3\] we solve for \(y\) using almost all algebra

OpenStudy (anonymous):

thank you for your patience :)

OpenStudy (anonymous):

a) subtract 3 get \[x-3=2\ln(\frac{y-1}{2})\]

OpenStudy (anonymous):

b) divide by 2, get \[\frac{x-3}{2}=\ln(\frac{y-1}{2})\]

OpenStudy (anonymous):

c) the only non algebra step write in equivalent exponential form , the base of the natural log is \(e\) so in exponential form this is \[\large e^{\frac{x-3}{2}}=\frac{y-1}{2}\]

OpenStudy (anonymous):

d) back to algebra multiply by 2 get \[\large 2e^{\frac{x-3}{2}}=y-1\]

OpenStudy (anonymous):

finally e) add 1 get \[\large 2e^{\frac{x-3}{2}}+1=y\] and that is your inverse

OpenStudy (anonymous):

domain x is an element of all the reals and y must be greater than 1?

OpenStudy (anonymous):

yes i would say that is right

OpenStudy (anonymous):

Can you help me with one more and I will leave you alone LOL

OpenStudy (anonymous):

since it is the reverse of the previous one with domain range switched

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i(x) = 1-3x10^2x+3 Domain and range, plus inverse :)

OpenStudy (anonymous):

i can't really read it

OpenStudy (anonymous):

looks like \[1-3x10^{2x}+3\]

OpenStudy (anonymous):

but maybe \[1-3\times 10^{2x+3}\]

OpenStudy (anonymous):

\[i(x)=1-3\times10^{2x+3}\]

OpenStudy (anonymous):

ok domain as usual is all real numbers range is \((-\infty,1)\) since \(10^x\) is never negative

OpenStudy (anonymous):

so range must be less than 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you want the inverse solve \[x=1-3\times 10^{2y+3}\] for \(y\) again almost all algebra

OpenStudy (anonymous):

when you get to \[-\frac{x-1}{3}=10^{2y+3}\] this time write in equivalent logarithmic form

OpenStudy (anonymous):

\[\log_{10} (-\frac{ y-1 }{ 3})=2x+3\]

OpenStudy (anonymous):

sorry x and y back to front